OpenStudy (anonymous):

somebody please help me..got test tonight but still dunn how to solve this... The base of a triangle increases at the rate 2cm per second, and height decreases at the rate of 1/2cm per second.Find the rate of change f its area when the base and height are of length 5cm... ? T.T helpp

OpenStudy (anonymous):

anybody ??? :(

OpenStudy (anonymous):

you know agent00smith ?

OpenStudy (agent0smith):

hmm.... we know \[A = 0.5 b h\] \[\frac{ db }{ dt } = 2 (cm/s)\] \[\frac{ dh }{ dt }= -0.5 (cm/s)\] We need \[\frac{ dA }{ dt } = ?\]

OpenStudy (anonymous):

chain rule ?

OpenStudy (anonymous):

but we have t differentiate A with respect to ??

OpenStudy (agent0smith):

Yeah that's what i'm trying to think of...

OpenStudy (anonymous):

o.O ok ... i really dunno how to relate that

zepdrix (zepdrix):

Hmm are we assuming this is a `right` triangle? :o

zepdrix (zepdrix):

I guess it would have to be, otherwise we need more information. :) Looks like agent was on the right track. Both \(\large b\) and \(\large h\) are variables. Just apply the product rule! :D

zepdrix (zepdrix):

Differentiate with respect to time*

OpenStudy (agent0smith):

Yeah that's what I was thinking... differentiate A wrt t

OpenStudy (anonymous):

huh ??? yeahh differentiate wrt time . that is wwhat we want to find

OpenStudy (anonymous):

im not clear with that

OpenStudy (agent0smith):

But then that seems to make the actual values of the base and height unnecessary, but maybe that's okay.... \[\frac{ dA }{ dt} = 0.5 \frac{ db }{ dt}\frac{ dh }{ dt}\]

zepdrix (zepdrix):

Woops, product rule! :O

zepdrix (zepdrix):

\[\large A=\frac{1}{2}bh\]\[\large A'=\frac{1}{2}b'h+\frac{1}{2}bh'\]Something like that, yes? :D

OpenStudy (anonymous):

huhh ????? how can...i dont undrstand...

OpenStudy (anonymous):

urmmm let me try that

OpenStudy (anonymous):

wow.....yeahh!!! correct answer !!! greatt !! thank you so much guys ! love uollsz!! :D

zepdrix (zepdrix):

yay team \:D/

OpenStudy (anonymous):

my mentors :)

OpenStudy (agent0smith):

ah yes, my bad